If only 7 percent kids can secure A grade in a paper, find the probability of atmost 2 out of 10 kids getting A grade in that paper?
1
Expert's answer
2020-01-31T09:45:37-0500

Let "p=0.07" and "q=1-p=0.93".

The probability of at most 2 out of 10 kids getting A grade in that paper is


"P(X\leq 2) = P(X=0) + P(X=1) + P(X=2) ="


"q^{10} + C_1^{10} q^{9} p^1+C_2^{10} q^{8} p^2="


"0.93^{10} +10*0.93^{9} * 0.07+45*0.93^{8} * 0.07^2 = 0.9717"